# Algebraic power series and their automatic complexity II: modulo prime powers

### v2

@@ 3. A numeration system for the automaton states
where the digits $T_{k}$ and $U_{k}$ belong to $D[x,y,y^{-1}]$. Then $T_{0}Q^{p^{\alpha-1}-1}\equiv U_{0}Q^{p^{\alpha-1}-1}\mod p$. Since $Q$ has a constant term which is nonzero modulo $p$, it is invertible, so $T_{0}\equiv U_{0}\mod p$, which implies $T_{0}=U_{0}$. Inductively, suppose that $T_{m}=U_{m}$ for $10\leq m\leq k-$ 1. Reducing Equation (4) modulo $p^{k+1}$ and expanding gives us
@@
$=\left(T_{0}+\dots+R_{k}\,(\tfrac{p}{Q})^{k}+(U_{k}Q+T_{k+1}^{\prime})\,(\tfrac{p}{Q})^{k+1}+\cdots+T_{\alpha-1}^{\prime}\,(\tfrac{p}{Q})^{\alpha-1}\right)Q^{p^{\alpha-1}-1},$
so that we can set $T_{k}=R_{k}$. ∎
@@
Let $P=xy^{2}+(x+1)y+x$. The coefficient sequence $a(n)_{n\geq 0}$ of the Furstenberg series $F$ satisfying $P(x,F)=0$ is $0,-1,1,-2,4,-9,21,-51,\dots$ and is a signed, shifted variant of the sequence of Motzkin numbers [20, A001006]. Let $p=2$ and $\alpha=3$. Consider the initial state $S_{0}\in\mathcal{R}_{8}[x,y,y^{-1}]$ of the automaton generating $(a(n)\bmod 8)_{n\geq 0}$. By definition,
@@
The set of Laurent polynomials in $\mathcal{R}_{p^{\alpha}}[x,y,y^{-1}]$ with a base-$\frac{p}{Q}$ representation is closed under $\lambda_{r,0}$ for $10\leq r\leq p-$ 1.
@@
For $m\in\{0,1,\dots,k\}$, the summand is 00 since $\Delta^{\beta}=0$. Therefore
@@
Multiplying both sides by $T$, we have $\frac{TQ(x^{p},y^{p})^{k+\beta}}{Q^{k+1}}=TZ$. Since the Laurent polynomial $Q(x^{p},y^{p})^{k+\beta}$ consists of terms with exponents that are multiples of $p$, applying $\Lambda_{r,0}$ to both sides gives $\Lambda_{r,0}(\frac{T}{Q^{k+1}})Q^{k+\beta}=\Lambda_{r,0}(TZ)$. In particular, $\Lambda_{r,0}(\frac{T}{Q^{k+1}})Q^{k+\beta}$ is a Laurent polynomial. Moreover, its $z$-degree and $y$-min-degree are as claimed, using the bounds in the previous paragraph on $\deg_{z}Z$ and $\operatorname{mindeg}_{y}Z$. Finally, since the coefficient of $x^{0}y^{-1}$ in $Q$ is 00, it follows from Lemma 9 that each nonzero monomial $c\,x^{I}y^{J}$ in $\Delta$ and $Z$ satisfies $J\geq-I$. Therefore if each nonzero monomial in $T$ also satisfies this constraint, then, again by Lemma 9, so does each nonzero monomial in $S$. ∎
@@
after substituting $j=m-k$. Switching the order of summation gives
@@
For $m=0$, the summand is 00 modulo $p^{k+1}$ since $\Delta$ is divisible by $p$. Therefore $\frac{Q^{k+1}}{P/y}\bmod p^{k+1}$ is a Laurent polynomial. Finally, the degree bounds follow from the fact that $\deg_{z}\Delta\leq\deg_{z}(P/y\bmod p^{k+1})$ for $z\in\{x,y\}$, and $\operatorname{mindeg}_{y}\Delta\geq\operatorname{mindeg}_{y}(P/y\bmod p^{k+1})$. ∎
@@ 4. First bounds on the size of the automaton
By Theorem 17, every state in $\mathcal{M}_{p^{\alpha}}$ has a base-$\frac{p}{Q}$ representation. In this section, we bound the digits of these representations, to give preliminary bounds on the size of $\mathcal{M}_{p^{\alpha}}$ in Corollaries 24 and 25. To do this, we will show that the base-$\frac{p}{Q}$ representations of most states live in the spaces $\mathcal{W}$ or $\mathcal{V}$ defined below in Equations (8) and (9).
@@
(8) $\mathcal{W}$ :=$\mathrel{\mathop{:}}=\{(T_{\alpha-1},\dots,T_{1},T_{0}):\text{$T_{k}\in W_{k}$ for each $k\in\{0,1,\dots,\alpha-1\}$}\}$
(9) $\mathcal{V}$ :=$\mathrel{\mathop{:}}=\{(T_{\alpha-1},\dots,T_{1},T_{0}):\text{$T_{k}\in V_{k}$ for each $k\in\{0,1,\dots,\alpha-1\}$}\}.$
Proposition 12 implies that $\mathcal{W}$ is a $D$-vector space, where the vector space addition operation is addition of the corresponding Laurent polynomials. The set of all tuples of the form $(0,\dots,0,x^{i}y^{j},0,\dots,0)$ with the appropriate bounds on $i$ and $j$ forms a basis of $\mathcal{W}$, in that every element of $\mathcal{W}$ is a unique $D$-linear combination of this basis. Figure 1 depicts the geometry of the monomials in $\mathcal{W}$ for a particular choice of values for $p$, $\alpha$, $h$, and $d$.
$(0,0)$$(\alpha-1,1-\alpha)=(2,-2)$$(\alpha h-1,1-\alpha)=(5,-2)$$(\alpha h-1,\alpha(d-1))=(5,9)$$(0,\alpha(d-1))=(0,9)$
@@ 5. Structure of the linear transformation $\lambda_{0,0}$
Proposition 26 below is analogous to [16, Proposition 13].
@@ 6. Period lengths of series expansions modulo $p$ and modulo $p^{\alpha}$
The following is a strengthening of Engstrom [9, Theorems 2 and 3], who bounds the period length of the coefficient sequence of a rational series $\frac{1}{R}$. In Theorem 29 we reduce his bound by a factor of $p$ when $e$ is a power of $p$.
@@
Lift $R$ and $R_{1},\dots,R_{k}$ to $\mathbb{Z}_{p}[z]$, and write $\frac{1}{R}=\sum_{n\geq 0}a(n)z^{n}\in\mathbb{Z}_{p}\llbracket z\rrbracket$. We follow Engstrom [9, Section 3.1]. For $i\in\{1,2,\dots,k\}$ and $j\in\{1,2,\dots,\deg R_{i}\}$, let $\rho_{i,j}$ be the $j$th root of $R_{i}(z)=0$. Then, for all $n\geq 0$, we have
@@
for some elements $c_{i,j,j^{\prime}}$ in the splitting field of $R$ over $\mathbb{Q}_{p}$. Engstrom [9, Lemma 4] implies that each $c_{i,j,j^{\prime}}$ is a $p$-adic algebraic integer. The appropriate generalization of Fermat’s little theorem implies that $\rho_{i,j}^{p^{\deg R_{i}}-1}\equiv 1\mod p$. Therefore the period length of $(\rho_{i,j}^{n}\bmod p)_{n\geq 0}$ divides $L$. It remains to bound the period lengths modulo $p$ of $\binom{n}{0},\dots,\binom{n}{e_{i}-1}$. A result of Ząbek [22, Théorème 3] implies that the period length of $(\binom{n}{j^{\prime}}\bmod p)_{n\geq 0}$ is $p^{\lfloor\log_{p}j^{\prime}\rfloor+1}\leq p^{\lfloor\log_{p}(e_{i}-1)\rfloor+1}\leq p^{\lfloor\log_{p}(e-1)\rfloor+1}=p^{\lceil\log_{p}e\rceil}$. The result follows. ∎
Theorem 30 (Engstrom [9, Theorem 8]).
@@
Let $T\in\mathcal{R}_{p^{\alpha}}[z]$ be a polynomial with $t\mathrel{\mathop{:}}=\deg T\geq 1$. If the coefficients of $z^{0}$ and $z^{t}$ in $T$ are nonzero modulo $p$, then the coefficient sequence of $\frac{1}{T}$ is periodic, and its period ends with exactly $t-1$ zeros.
@@
for all $i$ in the range $11\leq i\leq t-$ 1. Moreover, for all $n\geq 0$ we have the recurrence
@@
As $m\geq 1$, we can set $n=m-1$ in Equation (17). The periodicity of $a(n)_{n\geq 0}$ and Equation (16) with $i=1$ gives
@@
Since $c_{t}$ is invertible, we have $a(m-1)=0$. As $a(0)\neq 0$, this implies $m\geq 2$.
Therefore, we can set $n=m-2$ in Equation (17). Periodicity, $a(m-1)=0$, and Equation (16) with $i=2$ gives
@@
Finally, setting $n=m-t$ gives
@@
leaving exactly $t-1$ zeros at the end of the period of $a(n)_{n\geq 0}$. ∎
@@
Let $T\in\mathcal{R}_{p^{\alpha}}[z]$ such that $t\mathrel{\mathop{:}}=\deg T\geq 1$ and the coefficients of $z^{0}$ and $z^{t}$ in $T$ are nonzero modulo $p$. For all $i$ in the range $10\leq i\leq t-$ 1, the coefficient sequence of $\frac{1}{z^{-i}T}$ is periodic, and its period begins with $i$ zeros and ends with $t-1-i$ zeros.
@@ 7. Orbit size of a univariate Laurent polynomial under $\lambda_{0}$
In this section, our goal is to prove Corollary 43, which tells us that the orbit of a univariate Cartier operator $\lambda_{0}$ is finite, and which gives bounds on the transient and period length of this orbit.
We begin with the following result, which lets us restrict attention to Laurent polynomials $S$ with $\deg S\leq p^{\alpha-1}\deg(R\bmod p)$ in Theorems 38, 40, and 41. The proof uses the fact that if $f(x)=\left\lfloor\frac{x-p^{\alpha-1}r}{p}\right\rfloor+p^{\alpha-1}r$, then, for every $x\geq p^{\alpha-1}r$ and $n\geq\left\lfloor\log_{p}(x-p^{\alpha-1}r)\right\rfloor+1$, we have $f^{n}(x)=p^{\alpha-1}r$.
@@
As in Example 31, let $p=2$, $\alpha=2$, and $R=-z^{2}-z+1\in\mathcal{R}_{4}[z]$. Write $\frac{1}{R^{p^{\alpha-1}}}=\frac{1}{R^{2}}=\sum_{n\geq 0}a(n)z^{n}\in\mathcal{R}_{4}\llbracket z\rrbracket$. The sequence $a(n)_{n\geq 0}$ is periodic with period length $12$. In light of Lemma 35, we consider Laurent polynomials $S\in\mathcal{R}_{4}[z]$ such that $-1=1-p^{\alpha-1}\leq\operatorname{mindeg}S$ and $\deg S\leq p^{\alpha-1}\deg(R\bmod p)=4$. Let $j\in\{-1,0,1,2,3,4\}$, so that each monomial in $S$ is of the form $c\,z^{j}$. By Proposition 5,
@@
If $j=-1$, it can be verified that $\lambda_{0}(z^{-1})=0$. For $j\in\{0,1,2,3,4\}$, we show that $\Lambda_{0}^{4}(\frac{z^{j}}{R^{2}})=\Lambda_{0}^{2}(\frac{z^{j}}{R^{2}})$, so that
@@
Since $4n-j\equiv 16n-j\mod 12$ and $\left\lceil j/4\right\rceil=\left\lceil j/16\right\rceil$ for $j\in\{0,1,2,3,4\}$, it follows that $\Lambda_{0}^{4}(\frac{z^{j}}{R^{2}})=\Lambda_{0}^{2}(\frac{z^{j}}{R^{2}})$, as desired. By linearity, this implies $\lambda_{0}^{4}(S)=\lambda_{0}^{2}(S)$ for all $S\in\mathcal{R}_{4}[z]$ with $-1\leq\operatorname{mindeg}S$ and $\deg S\leq 4$.
@@
Engstrom’s Theorems 29 and 30 let us bound the eventual period length $m$ of $\frac{1}{R\bmod p}$, and in Corollary 34 we showed that the period length of the expansion of $\frac{1}{R^{p^{\alpha-1}}}$ divides $p^{2(\alpha-1)}m$. We will see that the orbit size under $\lambda_{0}$ is related to $m$. Since $\Lambda_{0}^{k}\!\left(\sum_{n\geq 0}a(n)z^{n}\right)=\sum_{n\geq 0}a(p^{k}n)z^{n}$, if $a(n)_{n\geq 0}$ has period length $m$ then $\operatorname{ord}_{m}(p)$ determines the generic orbit size under $\Lambda_{0}$. We will use the next lemma to evaluate it.
@@
The next several results establish the orbit size under $\lambda_{0}$ in various settings. We start with Theorem 38, where we assume that $\deg(R\bmod p)\geq 1$ and $\operatorname{mindeg}(R\bmod p)\in\{-1,0\}$. Then we consider $\operatorname{mindeg}(R\bmod p)\geq 1$ in Theorem 40. Finally in Theorems 41 and 42, we deal with the remaining possibilities. The statement of Corollary 43 covers all cases.
@@
If $j\leq 1$, then this sum is empty and therefore 00. So assume $j\in\{2,\dots,p^{\alpha-1}r\}$. It remains to show that $a(p^{t+\operatorname{ord}_{m}(p)}n-j)=0$ for all $n$ in the range of summation of the sum in Equation (20).
@@
Let $r\mathrel{\mathop{:}}=\deg(R\bmod p)$. The assumption that $r\in\{-1,0\}$ implies $e_{0}\in\{-1,0\}$. If $r=-1$, the only Laurent polynomial $S$ satisfying the constraints is $S=0$, and the conclusion holds.
@@
Let $j\in\{1-p^{\alpha-1},\dots,0\}$, so that $kz^{j}$ is a monomial in $S$. Equation (19) still holds, so that $\lambda_{0}^{n}(z^{j})=\Lambda_{0}^{n}\!\left(\frac{z^{j}}{R^{p^{\alpha-1}}}\right)R^{p^{\alpha-1}}$ for all $n\geq 0$. First assume $j=0$. The series $\frac{z^{j}}{R^{p^{\alpha-1}}}$ has transient length $1$. Applying $2(\alpha-1)$ iterations of $\Lambda_{0}$ produces a series with transient length at most $1$ and eventual period length dividing $m$. If $j\leq-1$ and $n\geq 2(\alpha-1)$, then $\Lambda_{0}^{n}\!\left(\frac{z^{j}}{R^{p^{\alpha-1}}}\right)$ is periodic with period length dividing $m$. Since $m$ divides $p-1$ in both cases, we have that the period length divides $p-1$, and therefore $p$ is congruent to $1$ modulo the period length. Hence further applications of $\Lambda_{0}$ leave the series fixed. This completes the case $b\neq 0$.
@@
We now have the following general result which includes Theorems 38, 40, 41, and 42.
@@
(21) $t=\max\!\left(\left\lfloor\log_{p}\max(e_{0},1)\right\rfloor+\alpha,\left\lceil\log_{p}\max(e_{1},\dots,e_{k},1)+2(\alpha-1)\right\rceil\right)$

### v3

@@ 3. A numeration system for the automaton states
where the digits $T_{k}$ and $U_{k}$ belong to $D[x,y,y^{-1}]$. Then $T_{0}Q^{p^{\alpha-1}-1}\equiv U_{0}Q^{p^{\alpha-1}-1}\mod p$. Since $Q$ has a constant term which is nonzero modulo $p$, it is invertible, so $T_{0}\equiv U_{0}\mod p$, which implies $T_{0}=U_{0}$. Inductively, suppose that $T_{m}=U_{m}$ for $0\leq m\leq k-1$. Reducing Equation (4) modulo $p^{k+1}$ and expanding gives us
@@
$=\left(T_{0}+\dots+R_{k}\,(\tfrac{p}{Q})^{k}+(U_{k}Q+T_{k+1}^{\prime})\,(\tfrac{p}{Q})^{k+1}+\dots+T_{\alpha-1}^{\prime}\,(\tfrac{p}{Q})^{\alpha-1}\right)Q^{p^{\alpha-1}-1},$
so that we can set $T_{k}=R_{k}$. For $k=\alpha-1$, the quotient $U_{\alpha-1}$ plays no role since $U_{\alpha-1}Q\,(\tfrac{p}{Q})^{\alpha}\equiv 0\mod p^{\alpha}$. ∎
@@
Let $P=xy^{2}+(x+1)y+x$. The coefficient sequence $a(n)_{n\geq 0}$ of the Furstenberg series $F$ satisfying $P(x,F)=0$ is $0,-1,1,-2,4,-9,21,-51,\dots$ and is a signed, shifted variant of the sequence of Motzkin numbers [22, A001006]. Let $p=2$ and $\alpha=3$. Consider the initial state $S_{0}\in\mathcal{R}_{8}[x,y,y^{-1}]$ of the automaton generating $(a(n)\bmod 8)_{n\geq 0}$. By definition,
@@
The set of Laurent polynomials in $\mathcal{R}_{p^{\alpha}}[x,y,y^{-1}]$ with a base-$\frac{p}{Q}$ representation is closed under $\lambda_{r,0}$ for $0\leq r\leq p-1$.
@@
For $m\in\{0,1,\dots,k\}$, the summand is 0 since $\Delta^{\beta}=0$. Therefore
@@
Multiplying both sides by $T$, we have $\frac{TQ(x^{p},y^{p})^{k+\beta}}{Q^{k+1}}=TZ$. Since the Laurent polynomial $Q(x^{p},y^{p})^{k+\beta}$ consists of terms with exponents that are multiples of $p$, applying $\Lambda_{r,0}$ to both sides gives $\Lambda_{r,0}(\frac{T}{Q^{k+1}})Q^{k+\beta}=\Lambda_{r,0}(TZ)$. In particular, $\Lambda_{r,0}(\frac{T}{Q^{k+1}})Q^{k+\beta}$ is a Laurent polynomial. Moreover, its $z$-degree and $y$-min-degree are as claimed, using the bounds in the previous paragraph on $\deg_{z}Z$ and $\operatorname{mindeg}_{y}Z$. Finally, since the coefficient of $x^{0}y^{-1}$ in $Q$ is 0, it follows from Lemma 9 that each nonzero monomial $c\,x^{I}y^{J}$ in $\Delta$ and $Z$ satisfies $J\geq-I$. Therefore if each nonzero monomial in $T$ also satisfies this constraint, then, again by Lemma 9, so does each nonzero monomial in $S$. ∎
@@
after substituting j=mkj=m-k. Switching the order of summation gives
@@
For $m=0$, the summand is 0 modulo $p^{k+1}$ since $\Delta$ is divisible by $p$. Therefore $\frac{Q^{k+1}}{P/y}\bmod p^{k+1}$ is a Laurent polynomial. Finally, the degree bounds follow from the fact that $\deg_{z}\Delta\leq\deg_{z}(P/y\bmod p^{k+1})$ for $z\in\{x,y\}$, and $\operatorname{mindeg}_{y}\Delta\geq\operatorname{mindeg}_{y}(P/y\bmod p^{k+1})$. ∎
@@
4. Compatibility of automata
Let $\mathcal{M}_{p^{\alpha}}$ be the automaton in Theorem 17 that generates the coefficients of $F\bmod p^{\alpha}$. In this section, we show that the base-$\frac{p}{Q}$ representations of states in $\mathcal{M}_{p^{\alpha+1}}$ project onto those of states in $\mathcal{M}_{p^{\alpha}}$. This implies that, when we vary $\alpha$, the automata $\mathcal{M}_{p^{\alpha}}$ support an inverse limit. In a previous article [20], the authors proved this for a sequence of automata reading most significant digit first instead, but no explicit description of its states was known. The advantage of the construction below is that it gives a description of the states as sequences of base-$\frac{p}{Q}$ digits. The results of this section are not needed for the remainder of the article. We begin with an example.
Example 19.
Notation.
Theorem 20.
Proof.
For $m=0$, we have $T_{0}^{\prime}=U_{0,0}\equiv\Lambda_{r,0}\!\left(T_{k}\hat{Q}^{-1}\right)\hat{Q}\mod p$. Therefore the 0th digits satisfy $\operatorname{dig}_{0}(\lambda_{r,0}^{(\alpha)}(S^{(\alpha)}))=\operatorname{dig}_{0}(\lambda_{r,0}^{(\beta)}(S^{(\beta)}))$.
Therefore
Corollary 21.
Proof.
[…]
It remains to show that for each $k\in\{0,1,\dots,\beta-1\}$, we have $\operatorname{dig}_{k}(S_{0}^{(\alpha)})=\operatorname{dig}_{k}(S_{0}^{(\beta)})$. From the proof of Theorem 17, the 0th digit of $S_{0}^{(\alpha)}$ is $T_{0}=\left(y\frac{\partial P}{\partial y}\bmod p\right)$, and this is independent of $\alpha$. Recursively, for $k\geq 1$, the $k$th digit of $S_{0}^{(\alpha)}$ is
@@ 5. First bounds on the size of the automaton
By Theorem 17, every state in $\mathcal{M}_{p^{\alpha}}$ has a base-$\frac{p}{Q}$ representation. In this section, we bound the digits of these representations, to give preliminary bounds on the size of $\mathcal{M}_{p^{\alpha}}$ in Corollaries 27 and 28. To do this, we will show that the base-$\frac{p}{Q}$ representations of most states live in the spaces $\mathcal{W}$ or $\mathcal{V}$ defined below in Equations (8) and (9).
@@
(8) $\mathcal{W}$ $\mathrel{\mathop{:}}=\{(T_{\alpha-1},\dots,T_{1},T_{0}):\text{$T_{k}\in W_{k}$ for each $k\in\{0,1,\dots,\alpha-1\}$}\}$
(9) $\mathcal{V}$ $\mathrel{\mathop{:}}=\{(T_{\alpha-1},\dots,T_{1},T_{0}):\text{$T_{k}\in V_{k}$ for each $k\in\{0,1,\dots,\alpha-1\}$}\}.$
The proof of Proposition 12 implies that $\mathcal{W}$ is a $D$-vector space, where the vector space addition operation is addition of the corresponding Laurent polynomials. The set of all tuples of the form $(0,\dots,0,x^{i}y^{j},0,\dots,0)$ with the appropriate bounds on $i$ and $j$ forms a basis of $\mathcal{W}$, in that every element of $\mathcal{W}$ is a unique $D$-linear combination of this basis. Figure 1 depicts the geometry of the monomials in $\mathcal{W}$ for a particular choice of values for $p$, $\alpha$, $h$, and $d$.
[…]
@@ 6. Structure of the linear transformation $\lambda_{0,0}$
Proposition 29 below is analogous to [18, Proposition 13].
@@ 7. Period lengths of series expansions modulo $p$ and modulo $p^{\alpha}$
The following is a strengthening of Engstrom [11, Theorems 2 and 3], who bounds the period length of the coefficient sequence of a rational series $\frac{1}{R}$. In Theorem 32 we reduce his bound by a factor of $p$ when $e$ is a power of $p$.
@@
Lift $R$ and $R_{1},\dots,R_{k}$ to $\mathbb{Z}_{p}[z]$, and write $\frac{1}{R}=\sum_{n\geq 0}a(n)z^{n}\in\mathbb{Z}_{p}\llbracket z\rrbracket$. We follow Engstrom [11, Section 3.1]. For $i\in\{1,2,\dots,k\}$ and $j\in\{1,2,\dots,\deg R_{i}\}$, let $\rho_{i,j}$ be the $j$th root of $R_{i}(z)=0$. Then, for all $n\geq 0$, we have
@@
for some elements $c_{i,j,j^{\prime}}$ in the splitting field of $R$ over $\mathbb{Q}_{p}$. Engstrom [11, Lemma 4] implies that each $c_{i,j,j^{\prime}}$ is a $p$-adic algebraic integer. The appropriate generalization of Fermat’s little theorem implies that $\rho_{i,j}^{p^{\deg R_{i}}-1}\equiv 1\mod p$. Therefore the period length of $(\rho_{i,j}^{n}\bmod p)_{n\geq 0}$ divides $L$. It remains to bound the period lengths modulo $p$ of $\binom{n}{0},\dots,\binom{n}{e_{i}-1}$. A result of Ząbek [24, Théorème 3] implies that the period length of $(\binom{n}{j^{\prime}}\bmod p)_{n\geq 0}$ is $p^{\lfloor\log_{p}j^{\prime}\rfloor+1}\leq p^{\lfloor\log_{p}(e_{i}-1)\rfloor+1}\leq p^{\lfloor\log_{p}(e-1)\rfloor+1}=p^{\lceil\log_{p}e\rceil}$. The result follows. ∎
Theorem 33 (Engstrom [11, Theorem 8]).
@@
Let $T\in\mathcal{R}_{p^{\alpha}}[z]$ be a polynomial with $t\mathrel{\mathop{:}}=\deg T\geq 1$. If the coefficients of $z^{0}$ and $z^{t}$ in $T$ are nonzero modulo $p$, then the coefficient sequence of $\frac{1}{T}$ is periodic, and its period ends with exactly t1t-1 zeros.
@@
for all $i$ in the range $1\leq i\leq t-1$. Moreover, for all $n\geq 0$ we have the recurrence
@@
As $m\geq 1$, we can set n=m1n=m-1 in Equation (17). The periodicity of $a(n)_{n\geq 0}$ and Equation (16) with $i=1$ gives
@@
Since $c_{t}$ is invertible, we have a(m1)=0a(m-1)=0. As $a(0)\neq 0$, this implies $m\geq 2$.
Therefore, we can set n=m2n=m-2 in Equation (17). Periodicity, a(m1)=0a(m-1)=0, and Equation (16) with $i=2$ gives
@@
Finally, setting n=mtn=m-t gives
@@
leaving exactly t1t-1 zeros at the end of the period of $a(n)_{n\geq 0}$. ∎
@@
Let $T\in\mathcal{R}_{p^{\alpha}}[z]$ such that $t\mathrel{\mathop{:}}=\deg T\geq 1$ and the coefficients of $z^{0}$ and $z^{t}$ in $T$ are nonzero modulo $p$. For all $i$ in the range $0\leq i\leq t-1$, the coefficient sequence of $\frac{1}{z^{-i}T}$ is periodic, and its period begins with $i$ zeros and ends with t1it-1-i zeros.
@@ 8. Orbit size of a univariate Laurent polynomial under $\lambda_{0}$
In this section, our goal is to prove Corollary 46, which tells us that the orbit under a univariate Cartier operator $\lambda_{0}$ is finite, and which gives bounds on the transient and period length of this orbit.
We begin with the following result, which lets us restrict attention to Laurent polynomials $S$ with $\deg S\leq p^{\alpha-1}\deg(R\bmod p)$ in Theorems 41, 43, and 44. The proof uses the fact that if $f(x)=\left\lfloor\frac{x-p^{\alpha-1}r}{p}\right\rfloor+p^{\alpha-1}r$, then, for every $x\geq p^{\alpha-1}r$ and $n\geq\left\lfloor\log_{p}(x-p^{\alpha-1}r)\right\rfloor+1$, we have $f^{n}(x)=p^{\alpha-1}r$.
@@
As in Example 34, let $p=2$, $\alpha=2$, and $R=-z^{2}-z+1\in\mathcal{R}_{4}[z]$. Write $\frac{1}{R^{p^{\alpha-1}}}=\frac{1}{R^{2}}=\sum_{n\geq 0}a(n)z^{n}\in\mathcal{R}_{4}\llbracket z\rrbracket$. The sequence $a(n)_{n\geq 0}$ is periodic with period length $12$. In light of Lemma 38, we consider Laurent polynomials $S\in\mathcal{R}_{4}[z]$ such that $1=1-p^{\alpha-1}\leq\operatorname{mindeg}S$ and $\deg S\leq p^{\alpha-1}\deg(R\bmod p)=4$. Let $j\in\{-1,0,1,2,3,4\}$, so that each monomial in $S$ is of the form $c\,z^{j}$. By Proposition 5,
@@
If j=1j=-1, it can be verified that $\lambda_{0}(z^{-1})=0$. For $j\in\{0,1,2,3,4\}$, we show that $\Lambda_{0}^{4}(\frac{z^{j}}{R^{2}})=\Lambda_{0}^{2}(\frac{z^{j}}{R^{2}})$, so that
@@
Since $4n-j\equiv 16n-j\mod 12$ and $\left\lceil j/4\right\rceil=\left\lceil j/16\right\rceil$ for $j\in\{0,1,2,3,4\}$, it follows that $\Lambda_{0}^{4}(\frac{z^{j}}{R^{2}})=\Lambda_{0}^{2}(\frac{z^{j}}{R^{2}})$, as desired. By linearity, this implies $\lambda_{0}^{4}(S)=\lambda_{0}^{2}(S)$ for all $S\in\mathcal{R}_{4}[z]$ with $1\leq\operatorname{mindeg}S$ and $\deg S\leq 4$.
@@
Engstrom’s Theorems 32 and 33 let us bound the eventual period length $m$ of $\frac{1}{R\bmod p}$, and in Corollary 37 we showed that the period length of the expansion of $\frac{1}{R^{p^{\alpha-1}}}$ divides $p^{2(\alpha-1)}m$. We will see that the orbit size under $\lambda_{0}$ is related to $m$. Since $\Lambda_{0}^{k}\!\left(\sum_{n\geq 0}a(n)z^{n}\right)=\sum_{n\geq 0}a(p^{k}n)z^{n}$, if $a(n)_{n\geq 0}$ has period length $m$ then $\operatorname{ord}_{m}(p)$ determines the generic orbit size under $\Lambda_{0}$. We will use the next lemma to evaluate it.
@@
The next several results establish the orbit size under $\lambda_{0}$ in various settings. We start with Theorem 41, where we assume that $\deg(R\bmod p)\geq 1$ and $\operatorname{mindeg}(R\bmod p)\in\{-1,0\}$. Then we consider $\operatorname{mindeg}(R\bmod p)\geq 1$ in Theorem 43. Finally in Theorems 44 and 45, we deal with the remaining possibilities. The statement of Corollary 46 covers all cases.
@@
If $j\leq 1$, then this sum is empty and therefore 0. So assume $j\in\{2,\dots,p^{\alpha-1}r\}$. It remains to show that $a(p^{t+\operatorname{ord}_{m}(p)}n-j)=0$ for all $n$ in the range of summation of the sum in Equation (20).
@@
Let $r\mathrel{\mathop{:}}=\deg(R\bmod p)$. The assumption that $r\in\{-1,0\}$ implies $e_{0}\in\{-1,0\}$. If r=1r=-1, the only Laurent polynomial $S$ satisfying the constraints is $S=0$, and the conclusion holds.
@@
Let $j\in\{1-p^{\alpha-1},\dots,0\}$, so that $kz^{j}$ is a monomial in $S$. Equation (19) still holds, so that $\lambda_{0}^{n}(z^{j})=\Lambda_{0}^{n}\!\left(\frac{z^{j}}{R^{p^{\alpha-1}}}\right)R^{p^{\alpha-1}}$ for all $n\geq 0$. First assume $j=0$. The series $\frac{z^{j}}{R^{p^{\alpha-1}}}$ has transient length $1$. Applying $2(\alpha-1)$ iterations of $\Lambda_{0}$ produces a series with transient length at most $1$ and eventual period length dividing $m$. If $j\leq-1$ and $n\geq 2(\alpha-1)$, then $\Lambda_{0}^{n}\!\left(\frac{z^{j}}{R^{p^{\alpha-1}}}\right)$ is periodic with period length dividing $m$. Since $m$ divides p1p-1 in both cases, we have that the period length divides p1p-1, and therefore $p$ is congruent to $1$ modulo the period length. Hence further applications of $\Lambda_{0}$ leave the series fixed. This completes the case $b\neq 0$.
@@
We now have the following general result by combining Theorems 41, 43, 44, and 45. Namely, Theorem 41 covers the case $\deg(R\bmod p)\geq 1$ and $\operatorname{mindeg}(R\bmod p)\in\{-1,0\}$, Theorem 43 covers $\operatorname{mindeg}(R\bmod p)\geq 1$, and Theorems 44 and 45 cover the remaining cases.
@@
(21) $t=\max\!\left(\left\lfloor\log_{p}\max(e_{0},1)\right\rfloor+\alpha,\left\lceil\log_{p}\max(e_{1},\dots,e_{k},1)\right\rceil+2(\alpha-1)\right)$